# Laser Interactions with Gas Jets: EMP Emission and Nozzle Damage

### v2

@@ VI.2 EMP Experiment on VEGA-3
A second experiment was conducted on VEGA-3 with a view to characterising the EMP and ions produced in laser-gas interactions. The laser energy before compressor was $\sim 30$ J and the pulse duration was 30 fs. Based on images of the focal spot taken at low energy, 21% of the energy on-target was contained within the first Airy disk[23] for an on-target intensity of order $\sim 10^{20}$ W cm-2 at best compression, similar to the intensity reported for the experiment in the previous section. The laser contrast was $10^{-12}$ at $0.1$ ns, with no significant pre-pulses[23]. A Prodyn RB-230(R) radiation-hardened B-dot probe was used to measure the amplitude of the magnetic field at different distances from the gas jet, while two diamond time of flight detectors and a Thomson parabola were used to monitor charged particle emission. Signal degradation due to EMP pick-up in the unshielded oscilloscope prevented the easy identification of a photo-peak in the time of flight data, so quantitative analysis of the ion spectrum is not possible. B-dot signals were recorded on a Rohde&Schwarz RTO64 digital oscilloscope with 2 GHz bandwidth and 10 GS/s sample rate. The B-dot probe was oriented so that its axis was parallel to the ground and sensitive to an azimuthal magnetic field relative to the nozzle axis.
@@
Both solid and gaseous targets were studied during the experiment. The gas was a 3:97 H2-He mix at $\sim 1000$ bar backing pressure, forced vertically downwards through the nozzle aperture to the laser focus. On shots with solid targets, the nozzle and solenoid valve assembly (Fig. 6) was replaced by a bracket and multi-foil array. The foils were made from $6\leavevmode\nobreak\ \mu$m-thick aluminium. Figure 7 shows typical EMP waveforms for solid and gaseous targets, demonstrating a reduction by a factor $2-3$ in the peak magnetic field for the same laser parameters when solid targets were switched to gas. This is consistent with the results of Kugland et al. [8] on the PHELIX laser with an Ar gas jet and tallies with our theoretical estimates. The ChoCoLaT2 code [63] evaluates the charge produced by the VEGA-3 laser interacting with $6\leavevmode\nobreak\ \mu$m-thick Al foils to be of order 100 nC, which is twice higher than the 42 nC estimated for a gas in Sec. VI.1.
@@
Figure 8: Variation of EMP maximum magnetic field with distance from the gas jet. Data was collected with the B-dot probe positioned at $\phi=60^{\circ}$ to the laser axis, with line of sight to the target occluded. The fitted curve is for a $3$ cm-tall antenna with the angle between the antenna axis and the probe assumed constant at $\theta=90^{\circ}$ for the different distances.
@@ VI.2.1 Comparison with EMP Emission Model
VII Conclusions
@@
We have explored various mechanisms of gas ionisation and nozzle damage in laser-gas interactions. Depending on the gas density and plasma temperature, an electrical discharge and EMP will be triggered either by plasma expansion from the laser focus or electrical breakdown of the ambient gas. The relatively small plasma volume produced in laser-gas interactions leads to a small plasma capacitance and this in turn produces a smaller ejected charge and lower level of EMP compared to solid targets. Nozzle damage is caused by plasma ions depositing energy in the nozzle surface rather than Ohmic heating from the discharge current. Two experiments on the VEGA-3 laser have provided useful supporting evidence for our model. The minimum speed of gas ionisation - inferred from optical interferometry - is consistent with collisional ionisation by ions in an expanding plasma or with the electric discharge.
@@ Acknowledgements.
The authors would like to thank the staff at the VEGA-3 Laser Facility for their hard work over the course of several experiments. This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No. 101052200 — EUROfusion). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. We acknowledge GENCI for providing us access to the Joliot-Curie supercomputer (grants 2021-A0130512993 and 2022-A0130512993). We also acknowledge Grant PID2021-125389OA-I00, funded by MCIN/ AEI / 10.13039/501100011033 / FEDER, UE and “ERDF A way of making Europe”, funded by the European Union. V. Ospina-Bohórquez and C. Vlachos acknowledge support from the LIGHT S&T Graduate Program (PIA3 Investment for the Future Program, ANR-17-EURE-0027). This work received funding from the European Union’s Horizon 2020 research and innovation program through the European IMPULSE project under grant agreement No 871161 and from LASERLAB-EUROPE V under grant agreement No 871124. It is published as part of the international project “PMW”, co-financed by the Polish Ministry of Science and Higher Education within the framework of the scientific financial resources for 2021-2022 under the contract no 5205/CELIA/2021/0 (project CNRS no 239915). Finally, we acknowledge the financial support of the IdEx University of Bordeaux / Grand Research Program “GPR LIGHT”.

### v3

@@ VI.2 EMP Experiment on VEGA-3
A second experiment was conducted on VEGA-3 with a view to characterising the EMP and ions produced in laser-gas interactions. The laser energy before compressor was $\sim 30$ J and the pulse duration was 30 fs. Based on images of the focal spot taken at low energy, 21% of the energy on-target was contained within the first Airy disk [23] for an on-target intensity of order $\sim 10^{20}$ W cm-2 at best compression, similar to the intensity reported for the experiment in the previous section. The laser contrast was $10^{-12}$ at $0.1$ ns, with no significant pre-pulses [23]. A Prodyn RB-230(R) radiation-hardened B-dot probe was used to measure the amplitude of the magnetic field at different distances from the gas jet, while two diamond time of flight detectors and three Thomson parabolas were used to monitor charged particle emission. Signal degradation due to EMP pick-up in the unshielded oscilloscope prevented the easy identification of a photo-peak in the time of flight data, so quantitative analysis of the ion spectrum is not possible. A proton spectrum from the Thomson parabola positioned at $90^{\circ}$ to the laser axis is shown in Fig. 3c. The data is integrated over five shots, with a lower proton sensitivity limit of $\sim 1.3$ MeV. The proton yield increases towards lower energies, with a maximum of $\sim 3\times 10^{11}$ MeV-1sr-1 at the low-energy limit. B-dot signals were recorded on a Rohde&Schwarz RTO64 digital oscilloscope with 2 GHz bandwidth and 10 GS/s sample rate. The B-dot probe was oriented so that its axis was parallel to the ground and sensitive to an azimuthal magnetic field relative to the nozzle axis.
@@
Both solid and gaseous targets were studied during the experiment. The gas was a 3:97 H2-He mix at $\sim 1000$ bar backing pressure, forced vertically downwards through the nozzle aperture to the laser focus. The electron density in the laser focal region was close to $10^{21}$ cm-3. On shots with solid targets, the nozzle and solenoid valve assembly (Fig. 8) was replaced by a bracket and multi-foil array. The foils were made from $6\leavevmode\nobreak\ \mu$m-thick aluminium. Figure 9 shows typical EMP waveforms for solid and gaseous targets, demonstrating a reduction by a factor $2-3$ in the peak magnetic field for the same laser parameters when solid targets were switched to gas. This is consistent with the results of Kugland et al. [8] on the PHELIX laser with an Ar gas jet and tallies with our theoretical estimates. The ChoCoLaT2 code [70] evaluates the charge produced by the VEGA-3 laser interacting with $6\leavevmode\nobreak\ \mu$m-thick Al foils to be of order 100 nC, which is twice higher than the 42 nC estimated for a gas in Sec. VI.1.
@@
Figure 10: Variation of EMP maximum magnetic field with distance from the gas jet. Data was collected with the B-dot probe positioned at $\phi=60^{\circ}$ to the laser axis, with the line of sight to the target occluded. The fitted curve is for a 3 cm-tall antenna with the angle between the antenna axis and the probe assumed constant at $\theta=90^{\circ}$ for the different distances.
@@ VI.2.1 Comparison with EMP Emission Model
VII Discussion
We examine here how EMP amplitude and nozzle damage vary on different laser systems. In Fig. 11, we consider the EMP produced by three different types of laser-gas interaction: (a) experiments at relatively low laser energy with under-dense gases, (b) experiments with PW-scale lasers and near-critical density gases, (c) experiments with high energy, longer pulse duration lasers and under-dense gas targets.
Figure 11: Variation of the EMP electric field located 1 m from a 3 cm-tall nozzle for different values of the total laser energy and pulse duration, calculated using the model from Sec. II. Laser and gas parameters have been chosen so that they are representative of experiments with: (a) under-dense gases on low-energy systems like the Gemini Target Area 2 laser [71], (b) near-critical gases and PW-scale lasers, such as VEGA-3, (c) under-dense gases and high energy, longer pulse duration lasers such as LMJ PETAL[72] and NIF ARC[73].
For all the calculations shown in Fig. 11, we assume 40% of the total laser energy shown on the $x$-axis is contained within the focal spot and that all the laser energy in the focus is absorbed over the full length of the plasma channel. The gas is He, ejected from a 3 cm-tall conducting nozzle, and the laser-to-hot-electron conversion efficiency is assumed to be 20% in panel (a), 30% in panel (b) and 40% in panel (c). For the relatively low-energy laser systems in panel (a), we assume a plasma channel radius of $r_{p}=20\,\mu$m, channel length $l_{p}=3$ mm and a gas pressure of $p_{g}=0.2$ bar, with the laser focused $h_{p}=3$ mm above the nozzle. In panel (b), we take $r_{p}=20\,\mu$m, $l_{p}=0.5$ mm, $p_{g}=10$ bar and $h_{p}=0.5$ mm. Then for the PETAL-type lasers in panel (c), we consider $r_{p}=40\,\mu$m, $l_{p}=1$ cm, $p_{g}=0.2$ bar and $h_{p}=3$ mm.
Calculations suggest that low-energy lasers can eject a few tens of nC from a plasma channel and produce EMP fields of a few kV/m at 1 m from the target. PW-class lasers are more disruptive, producing tens to hundreds of nC plasma charge and EMP fields of several tens of kV/m at 1 m from the gas. The strongest fields are expected for PETAL-class lasers, which can displace hundreds of nC to $\mu$C of charge and produce EMP fields of several hundred kV/m at metre-scale distances. These estimates agree well with recent measurements by W. Cayzac et al. at PETAL [74], which confirm that the EMP fields are generally a factor few lower than you would expect for solid target interactions.
The magnitude of the EMP can be controlled either by disrupting the discharge current as it propagates down the gas nozzle or through its relation to the amount of charge that escapes the plasma. A higher laser energy increases the energy available for conversion into hot electrons, while a higher intensity will increase the hot electron temperature. Increasing the laser energy and intensity will therefore tend to increase the escaping electron charge. A larger plasma volume or higher gas density will spread the laser energy over more particles and generally lower the average particle energy. A larger volume will also increase the plasma capacitance and reduce the electrostatic barrier potential for the escaping electrons.
Our model can also be used to estimate when a nozzle is likely to be damaged on a given laser system. By way of example, Fig. 12 shows the damage produced by a VEGA-3 type laser focused at different heights, $h_{p}$, above a metallic nozzle. The gas pressure, $p_{g}$, is varied up to $10$ bar and the different colours correspond to the ion energy deposited in the nozzle as a fraction of the energy required to melt the nozzle ($\sim 1.3$ kJ/g). Any contour value above $1.0$ therefore implies that the nozzle will be damaged. An artificial, exponentially decreasing density profile with $500\,\mu$m scale length is assumed, while the other parameters are as given in Table 3 for the VEGA-3 experiment. Damage occurs for laser-nozzle distances below a few hundred microns and peak gas pressures above a few bar. Damage is most easily avoided by focusing the laser further from the nozzle, where the gas density is lower and the ion flux on the nozzle will be reduced.
Figure 12: Nozzle damage factor (ratio of the plasma ion energy deposited in the nozzle per mass of heated material divided by the nozzle melt threshold) for different values of the laser-nozzle distance and gas pressure, calculated using the model from Sec. II. An exponentially decaying gas density profile with $500$ $\mu$m scale length is assumed, where the peak He gas density is taken at the nozzle surface ($h_{p}=0$). A bold white line marks where the nozzle damage factor is equal to 1, corresponding to a deposited energy-per-mass equal to the melt threshold of Cu ($\sim 1.3$ kJ/g).
It is important to bear in mind some of the limitations of our model. Many of the physical quantities used as inputs depend on each other in complicated ways, such as the relationship between the plasma channel volume and laser focal spot size, or between gas pressure and the hot electron conversion efficiency. Without additional information about these connections, the model will not be able to accurately predict how the EMP fields scale with physical parameters. Our model also does not account for plasma acceleration processes like shock [75] or laser wakefield [76] acceleration, which can affect the electron and ion distributions in complicated ways.
VIII Conclusions
@@
We have explored various mechanisms of gas ionisation and nozzle damage in laser-gas interactions. Depending on the gas density and plasma temperature, an electrical discharge and EMP can be triggered either by plasma expansion from the laser focus or electrical breakdown of the ambient gas. The relatively small plasma volume produced in laser-gas interactions leads to a small plasma capacitance and this in turn produces a smaller ejected charge and lower level of EMP compared to solid targets. Nozzle damage is caused by plasma ions depositing energy in the nozzle surface rather than Ohmic heating from the discharge current. Two experiments on the VEGA-3 laser have provided useful supporting evidence for our model. The minimum speed of gas ionisation - inferred from optical interferometry - is consistent with collisional ionisation by ions in an expanding plasma or with the electric discharge.
@@ Acknowledgements.
The authors thank the staff at the VEGA-3 Laser Facility for their hard work over the course of several experiments. This work has been carried out within the framework of the EUROfusion Consortium, funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No. 101052200 — EUROfusion). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. We acknowledge GENCI for providing us access to the Joliot-Curie supercomputer (grants 2021-A0130512993 and 2022-A0130512993). We also acknowledge Grant PID2021-125389OA-I00, funded by MCIN/ AEI / 10.13039/501100011033 / FEDER, UE and “ERDF A way of making Europe”, funded by the European Union. V. Ospina-Bohórquez and C. Vlachos acknowledge support from the LIGHT S&T Graduate Program (PIA3 Investment for the Future Program, ANR-17-EURE-0027). This work received funding from the European Union’s Horizon 2020 research and innovation program through the European IMPULSE project under grant agreement No 871161 and from LASERLAB-EUROPE V under grant agreement No 871124. It is published as part of the international project “PMW”, co-financed by the Polish Ministry of Science and Higher Education within the framework of the scientific financial resources for 2021-2022 under the contract no 5205/CELIA/2021/0 (project CNRS no 239915). Finally, we acknowledge the financial support of the IdEx University of Bordeaux / Grand Research Program “GPR LIGHT”.